AZ AO science

Juan Carlos left the unit powered up after completing daytime operations.

3:45 Cooling OCAM in preparation for ESM target. Seeing much better than predicted to we will work to execute a high priority ESM target then likeyly come back to ESM or AO during the second half. <1". (0.68" on the DIMM)

4:49 On our first ESM target. Catalog magnitude 14.26. Preset sent with 15.2 as the magnitude to force bin 4, frequency 400. Estimagted magnitude on teh WFS is 15.7. Seeing is 0.79" on the DIMM. Tried adjusting freq both up and down. HIgher frequency actually made it worse wo we have

Tried to save some tracking numbers but it didn't save. We did not meet this PI's crazy expectations so we moved on. MOVing to seeing limited

8:53 ESM preset sent. Railing the camera lens. This is a MOS mask so our rotator position is not flexible. Seeing has deteriorated so we may not be able to get this target anyways. Bin4, freq 400, mod 3, emgain 600.  Rotator at 40deg, pupil diff ~0.22. 

This ESM target was executed 20210405 and we observed for 3 hours. Tried again to save optical loop data but no luck. I finally got an error when the shell rested during a reconfig. Timeout.

Seeing deteriorated and was highly variable. We tested to see if pointed away from teh wind would help and if we could execute the AO program but no luck tonight. We abandoned ESM/AO for the night at 9:18

9:30 Warmed OCAM

Sorry for sparse log - helping on science camera while SciCam operator experience network issues.

-- JenniferPower - 08 Apr 2021
Topic revision: r3 - 08 Apr 2021, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback